Multiply
I am not sure how to do this problem can you please help me!
Thank you

Multiply
(x-3/2) (x+2/3)

Answer by jessica43(140)

For this problem you are going to use distribution in multiplication: multiply the 1st number by the 3rd and 4th, add them, then multiply the 2nd number by the 3rd and 4th, add them.
(x-(3/2))*(x+(2/3))=
(x)*(x) + (x)*(2/3) + (-3/2)*(x) + (-3/2)*(2/3)=
x^2 + (2/3)x - (3/2)x - 1=
Now you combine like terms to simplify. In this case, (2/3)x and (-3/2)x are like terms, so you add the coefficents to get:
x^2 -(5/6)x -1
So your final answer is x^2 - (5/6)x - 1.